Jacob is 21 years younger than twice marion's age. Jacob is 13. How old is marion?
Answer by ankor@dixie-net.com(22740)

let j = Jacob's present age
let m = Marion's
:
Write an equation for each statement
:
Jacob is 21 years younger than twice Marion's age.
j = 2m - 21
Jacob is 13.
j = 13
How old is Marion?
replace j with 13 in the 1st equation
13 = 2m - 21
add 21 to both sides
13 + 21 = 2m
34 = 2m
divide both sides by 2
m = 34/2
m = 17 yrs is Marion's age
:
:
:
Check the solution in the 1st statement
" Jacob is 21 years younger than twice Marion's age. "
13 = 2(17) - 21
13 = 34 - 21